OpenStudy (anonymous):

I need help with converting y-235x = 4137 to point slope form y - y1 = m(x - x1). Can someone help me?

OpenStudy (anonymous):

hi on this equation what is your slope?

OpenStudy (anonymous):

y-235x = 4137 adding 235x to both sides we get?:

OpenStudy (anonymous):

yessssss correct

OpenStudy (anonymous):

just do the pattern in the formula point slope form y - y1 = m(x - x1). y =235x-4137 do some factoring here

OpenStudy (anonymous):

y - y1 = m(x - x1). y - 0 = 235(x - x1). find x1 using 235 and 4137 using division or factoring

OpenStudy (anonymous):

to find x1 235 x1=4137, what is x1?

zepdrix (zepdrix):

I'm not sure if this is really necessary, but if you split the 4137 into 2 numbers, you can find a nice multiple of 235, so it will make your division a little bit easier. \[\large y=235x-4137\]\[\large y=235x-3995-142\]\[\large y+142=235x-3995\]Then do what mark is suggesting, factor out the 235 from each term on the right :D \[\large y+142=235(x-17)\] Just an idea! :) Looks like you're on the right track already though ^^ so don't mind me.

OpenStudy (anonymous):

yes plug them in to the formula y - y1 = m(x - x1). y - 0 = 235(x - x1).

OpenStudy (anonymous):

ok zepdrix got it... :D

OpenStudy (anonymous):

4137 / 235 = 17.60425532 on this one we will have an excess of 0.60425532

OpenStudy (anonymous):

you got it now:D

zepdrix (zepdrix):

You don't have to split up the numbers if you don't want ^^ it makes it look nicer, but i don't suppose it's entirely necessary. Just leave your x1 term as a fraction like this. \[\large y-0=235(x-x1)\]\[\large y-0=235(x-\frac{4137}{235})\] If splitting up the 4137 into a number divisible by 235 is confusing you, then just do it this way :D

OpenStudy (anonymous):

as long as you know where did you get 142 ok? :D

OpenStudy (anonymous):

ok have fun now and good luck .. :D